OpenStudy (anonymous):

Help please! Please please please!(: With the problem below, Determine the following: g(2)=____ g(4)=____ absolute max of g(x) is when x is ____ and is the value ____.

OpenStudy (anonymous):

Let: \[f(x)=0 , x<4 \] \[f(x)=5 , -4\le x <1\] \[f(x)=-2 , 1\le x <3\] \[f(x)=0 , x \ge3\] and \[g(x)=\int\limits_{-4}^{x} f(t) dt\]

OpenStudy (anonymous):

really would help to draw a picture, which is just a bunch of horizontal lines

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

the integral is the area of the rectangle so \[\int_{-4}^1f(x)dx=4\times 5=20\] which is the largest it can be, because the stuff after than it negative (below the \(x\) axis)

OpenStudy (anonymous):

To find the extrema of \(g(x)\), apply the first derivative test. If \(\displaystyle g(x)=\int_{-4}^xf(t)~dt\), then \(g'(x)=f(x)\). Find your critical points (i.e. when \(f(x)=0\)), which shouldn't be too hard, then determine the intervals on which \(g\) is increasing/decreasing.

OpenStudy (anonymous):

therefore the max is \(20\) when \(x=1\)

OpenStudy (anonymous):

this max is not 20

OpenStudy (anonymous):

hmm i think the critical points are not where the derivative is 0, but rather where the derivative is undefined small matter though

OpenStudy (anonymous):

oh of course, sorry, the max is \(5\times 5=25\) my mistake

OpenStudy (anonymous):

@satellite73, both, actually: http://en.wikipedia.org/wiki/Critical_point_(mathematics) (first sentence)

OpenStudy (anonymous):

@SithsAndGiggles i mean in this example

OpenStudy (anonymous):

Right, I was thinking more generally :P

OpenStudy (anonymous):

I'm so confused. so how is it determined at g(2) and g(4)?

OpenStudy (anonymous):

in general of course critical point is where derivative is zero or undefined in this example the important critical point would be at the change in the definition of the function, where the derivative would not be defined

OpenStudy (anonymous):

take the area of the rectangle at \[g(2)=\int_{-4}^2f(x)dx\] you have \(25\) for the area of the rectangle with base 5 and height 5, then \(-2\) for the area of the rectangle with base 1 and "height" \(-2\)

OpenStudy (anonymous):

therefore \(g(2)=23\)

OpenStudy (anonymous):

so with that g(4)=21 right?

OpenStudy (anonymous):

sorry i got closed out

OpenStudy (anonymous):

yea, \(g(4)=21\) right

OpenStudy (anonymous):

yay! thanks!!(: